OpenStudy (anonymous):

An equation is shown below: 8(2x – 14) + 13 = 4x – 27 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points) Please help i will fan????????!!!!!!!!!!!

OpenStudy (anonymous):

Im in FLVS

OpenStudy (sohailiftikhar):

8(2x-14)+13=4x-27 now multiply 8 inside you will get 16x-112+13=4x-27 now combine the x terms 16x-4x=112-40 12x=72 x=6

OpenStudy (campbell_st):

subtract 13 from both sides 8(2x -14)= 4x - 40 now divide both sides by 4 2(2x -14) = x - 10 now it should be easier to handle

OpenStudy (anonymous):

First do the distributive property. Then combine like terms. Subtract both sides of the equation by whatever value you are using. Now divide both sides. (I wont tell u the answer.) Now you get x alone.